OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2

OpenStudy (radar):

\[\sqrt{2}\times \sqrt[3]{2}=2^{(1/2 + 1/3)}\]

OpenStudy (anonymous):

\[2^{\frac{ 5 }{ 6 }}\]